# Real exponential sums over primes and prime gaps

### v2

@@ 3. Some technical lemmas
Lemma 3.8.
The function $\Gamma(z)\Omega((1-\lambda)(z+1))$ has exponential decay along vetical lines in the half-plane $\Re(z)>-1$.
Proof.
If $\Re(z)>-1$, then
$\Re((1-\lambda)(z+1))=(1-\lambda)[\Re(z)+1]>0,$
so that
$\Omega((1-\lambda)(z+1))=\Phi(1+(1-\lambda)(z+1))-\displaystyle\frac{1}{(1-\lambda)(z+1)}=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+(1-\lambda)(z+1)}}-\displaystyle\frac{1}{(1-\lambda)(z+1)}.$
Thus, we just need to find an upper bound for the sum $\sum_{p}\displaystyle\frac{\log(p)}{p^{1+(1-\lambda)(z+1)}}$. In fact,
$\left|\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+(1-\lambda)(z+1)}}\right|\leq\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+(1-\lambda)[\Re(z)+1]}}=O\left(\left(\displaystyle\frac{1}{(1-\lambda)[\Re(z)+1]}\right)^{2}\right).$
@@ 4. Final thoughts
Now, let’s make three comments about result 2.18. Firstly, an old problem from Erdős asks (see [4]) if the sequence
@@
Obviously, the proof is the same as the previous one. Is Sierpiński’s conjecture true for every $n\geq 2$?
@@
This function is already good enough to make all the steps described in section 2 work (including steps $8$ and $9$ of Newman’s proof), except for step $10$, in which its respective function $\Psi$ starts to be written using other functions, which complicates things a lot. The solution to this technicality was then to try to find a function $w$ whose associated cumulative function was asymptotic to $\exp(cx^{1-\lambda})$; that’s the case of
@@
I was amazed at how well this function behaves in relation to the necessary adaptations proposed here (in particular, by step 3, whose change of variables is explicit), but then I got stuck looking for how to realize step 8. After many unsuccessful attempts, I thought about expanding $\Psi(s)$ around $\lambda=1$, as follows: first, write
$\Psi(s)=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{\lambda}\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+s}}\cdot\displaystyle\frac{p^{s+1-\lambda}}{\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}$
Now, expanding the rightmost term around $\lambda=1$ for each prime $p$, one gets
(4.1) $\Psi(s)=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+s}}\cdot\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(1-\lambda)^{j}\log(p)^{j}f_{j}(s\log(p))}{j!},$
in which these $\{f_{j}(x)\}_{j\geq 0}$ are a sequence of polynomials in $\mathbb{Q}[x]$ defined recursively by
$f_{0}(x)=1$
and
$f_{j+1}(x)=f_{j}(x)-x\displaystyle\sum_{k=0}^{j}\displaystyle\binom{j}{k}\displaystyle\frac{f_{j-k}(x)}{k+2},$
for all $j\geq 0$, additionally having the marvelous property that the function
$s\mapsto\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{s+1}},$
initially defined for $\Re(s)>0$, extends analytically to $\Re(s)\geq 0$, for all $j\geq 1$. This makes us very willing to change the order of the sums in equation 4.1. Indeed, if we could do that, we would have
$\begin{split}\Psi(s)&=\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{1+s}}\\ &=\Phi(1+s)+\displaystyle\sum_{j=1}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{1+s}},\end{split}$
in such a way that by subtracting $\frac{1}{s}$ from each side of this equation, we would have
$\Psi(s)-\displaystyle\frac{1}{s}=\Omega(s)+\displaystyle\sum_{j=1}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{1+s}},$
thereby proving theorem 2.18. But I was unable to show that we could exchange the order of the sums.
@@ Acknowledgments
The author would like to thank his colleagues of IFSP, campus Itaquaquecetuba, for the possibility of obtaining the leave for the accomplishment of the postdoctoral research at IME - USP, and to his supervisor Hugo Luiz Mariano for all the support given during this period. He also thanks Professor Kai (Steve) Fan for pointing out an inconsistency in the first draft of the article (in the proof of proposition 2.8), and Professor Carl Pomerance for his attention to this matter.

### v3

@@ 4. Final thoughts
Now, let’s make three comments about result 2.22. Firstly, an old problem from Erdős asks (see [4]) if the sequence
@@
Clearly, the proof is the same as the previous one. Is Sierpiński’s conjecture true for every $n\geq 2$?
@@
This function is already good enough to make all the steps described in section 2 work (including steps $8$ and $9$ of Newman’s proof), except for step $10$, in which its respective function $\Psi$ starts to be written using other functions, which complicates things a lot. The solution to this technicality was then to try to find a function $w$ whose associated cumulative function had an asymptotic behavior similar to the function $\exp(cx^{1-\lambda})$; that’s the case of
@@
I was amazed at how well this function behaves in relation to the necessary adaptations proposed here (in particular, by step 3, whose change of variables is explicit), but then I got stuck looking for how to realize step 8. After many unsuccessful attempts, I thought about expanding $\Psi(s)$ around $\lambda=1$, as as described in steps $8$ and $9$, but initially I wasn’t sure about the possibility of changing the order of the double sum. After realizing that the double sum converges locally uniformly, I noticed that we can actually change the order of the double sum, and the result came out.
Other approaches I tried to finalize the proof of formulas 1.1 and 1.2 were to try using the Mellin transform (as per version $2$ of the article on ArXiv), or to make small modifications to the function $\Psi(s)$, in order to try to make it more tangible, as follows: first, I defined the functions
Obviously these functions are analytic on $\Re(s)>0$. Then, I proved the following equivalences:
Proposition 4.10.
The following statements are equivalent:
  1. i)
    The function $\Psi(s)-1/s$, initially defined for $\Re(s)>0$, extends analytically to $\Re(s)\geq 0$.
  2. ii)
    The function $\Xi(s)-1/s$, initially defined for $\Re(s)>0$, extends analytically to $\Re(s)\geq 0$.
  3. iii)
    The function $\tau(s)-1/s$, initially defined for $\Re(s)>0$, extends analytically to $\Re(s)\geq 0$.
Proof.
i) $\Rightarrow$ ii) By assumption, there is an analytic function $s\mapsto F(s)$, defined on an open set containing $\Re(s)\geq 0$, such that
$\Psi(s)-\displaystyle\frac{1}{s}=F(s),$
for all $s\in\mathbb{C}$ such that $\Re(s)>0$, that is
$\Xi(s)-\displaystyle\frac{1}{s}=(1-\lambda)F((1-\lambda)s),$
ii) $\Rightarrow$ i) By assumption, there is an analytic function $s\mapsto G(s)$, defined on an open set containing $\Re(s)\geq 0$, such that
$\Xi(s)-\displaystyle\frac{1}{s}=G(s),$
for all $s\in\mathbb{C}$ such that $\Re(s)>0$, that is,
for all $s\in\mathbb{C}$ such that $\Re(s)>0$. This shows that $\Psi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$.
ii) $\Leftrightarrow$ iii) Note that for all $s\in\mathbb{C}$ such that $\Re(s)>0$ holds
$\begin{split}\Xi(s)-\displaystyle\frac{1}{s}&=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{\lambda}\exp[s(p^{1-\lambda}-1)]}-\displaystyle\frac{1}{s}=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{\lambda}\exp(sp^{1-\lambda})\exp(-s)}-\displaystyle\frac{1}{s}\\ &=\exp(s)\tau(s)-\displaystyle\frac{1}{s}=\exp(s)\left[\tau(s)-\displaystyle\frac{1}{s}-\left(\displaystyle\frac{1}{s\exp(s)}-\displaystyle\frac{1}{s}\right)\right].\end{split}$
@@ Acknowledgments
The author would like to thank his colleagues of IFSP, campus Itaquaquecetuba, for the possibility of obtaining the leave for the accomplishment of the postdoctoral research at IME - USP, and to his supervisor Hugo Luiz Mariano for all the support given during this period. He also thanks Professor Kai (Steve) Fan for pointing out an inconsistency in the first draft of the article (in the proof of proposition 2.8), and Professor Carl Pomerance for his attention to this matter. Additionally, he also thanks Professor Gregory Debruyne for pointing out a flaw in the use of Mellin’s inverse transform in version $2$ of this article.